OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m? A)x^2 + y^2 - 2ax - 2by + (a^2 + b^2 - m^2) = 0 B)x^2 + y^2 + 2ax + 2by + (a2 + b^2 - m^2) = 0 C)x^2 + y^2 - 2ax - 2by + (a + b - m^2) = 0 D)x^2 + y^2 + 2ax + 2by + a^2 + b^2 = -m^2

OpenStudy (mrnood):

all you need to do for this is take the more familiar form (x-a)^2 + (y-b)^2 = m^2 expand the brackets as follows (x-a)(x-a)=? (y-b)(y-b) = then rearrange the equation so it looks like one of your answers

OpenStudy (anonymous):

ok so how would i do that?

OpenStudy (mrnood):

I have given you the mains steps already. in th efirst equation you have (x-a)^2 That is the same as (x-a)(x-a) First write what that multiplies out to (use FOIL?)

OpenStudy (anonymous):

okay so (x-a)^2 + (y-b)^2 = m^2 is the same as x^2-2ax+a^2 + y^2-2by+b^2-m^2 = 0 right? so the answer would be A?

OpenStudy (mrnood):

Well done - that's correct.